### v1

# A Principle of Targeted Intervention for Multi-Agent Reinforcement Learning

Abstract
Steering cooperative multi-agent reinforcement learning (MARL) towards desired outcomes is challenging, particularly when the global guidance from a human on the whole multi-agent system is impractical in a large-scale MARL. On the other hand, designing mechanisms to coordinate agents most relies on empirical studies, lacking a easy-to-use research tool. In this work, we employ multi-agent influence diagrams (MAIDs) as a graphical framework to address the above issues. First, we introduce interaction paradigms that leverage MAIDs to analyze and visualize existing approaches in MARL. Then, we design a new interaction paradigm based on MAIDs, referred to as targeted intervention that is applied to only a single targeted agent, so the problem of global guidance can be mitigated. In our implementation, we introduce a causal inference technique—referred to as Pre-Strategy Intervention (PSI)—to realize the targeted intervention paradigm. Since MAIDs can be regarded as a special class of causal diagrams, a composite desired outcome that integrates the primary task goal and an additional desired outcome can be achieved by maximizing the corresponding causal effect through the PSI. Moreover, the bundled relevance graph analysis of MAIDs provides a tool to identify whether an MARL learning paradigm is workable under the design of an interaction paradigm. In experiments, we demonstrate the effectiveness of our proposed targeted intervention, and verify the result of relevance graph analysis.
1 Introduction
Multi-agent reinforcement learning (MARL) provides a powerful framework for sequential decision-making in complex interacting systems, applied in autonomous systems including robotics hernandez2019survey ; panait2005cooperative ; cao2012overview . A central challenge in MARL is enabling effective coordination among agents that must operate with partial information in dynamic environments. Achieving robust coordination is often hindered by inherent difficulties such as non-stationarity papoudakis2019dealing , where concurrent learning destabilizes individual agent’s perspectives, making stable learning difficult. Furthermore, effective coordination often requires guiding a multi-agent system towards a desired outcome—an outcome with specific beneficial properties that standard MARL approaches may struggle to achieve consistently—especially when only maximizing a single, shared team reward (task goal) agogino2008analyzing ; huh2023multi .
To address these MARL challenges, researchers often employ guidance under various interaction paradigms to enhance performance, efficiency and applicability. Intrinsic rewards, for instance, can improve sample efficiency and direct agent behaviour by encouraging exploration or coordination jaques2019social ; du2019liir ; mohamed2015variational . Another approach incorporating human feedback, leverages human expertise to accelerate learning. This helps agents tackle objectives that are complex to specify via simple reward functions, and aligns agent behaviour to desired norms or human values christiano2017deep ; hu2022human ; mnih2015human .
All these interaction paradigms cannot be simply categorized by the conventional MARL learning paradigms (i.e., independent learning and centralized training).
While beneficial, applying global guidance to the entire multi-agent system in MARL is often impractical due to complexity, costs liang2024learning and safety concerns garg2024learning . Consider, for example, autonomous vehicles navigating a complex intersection or merging onto a highway. Ensuring safe and efficient passage requires intricate coordination. Also, providing simultaneous, specific feedback or instructions to all vehicles is infeasible, due to challenges such as complex safety validation and lack of universal communication protocols taeihagh2019governing . Yet, effective group coordination remains essential. An intriguing possibility arises if guidance is provided to a single targeted vehicle, e.g., instructing it to adjust its speed or yield at a critical moment. The resulting behaviour can serve as a pattern, enabling surrounding vehicles to coordinate. This scenario highlights a critical challenge where global guidance is essential but prohibitively difficult or costly, leading directly to the central question motivating our work:
Can effective coordination still be achieved by assigning an additional desired outcome to a single agent, relying on its influence over the rest of the agents in a multi-agent system?
Addressing the challenge of designing effective targeted intervention requires a principled framework to model and analyze how guiding a single agent impacts the overall multi-agent system. To this end, we propose leveraging multi-agent influence diagrams (MAIDs) koller2003multi , as a formal graphical language to encode complex strategic dependencies and information flow within the MARL learning process. Crucially, their bundled relevance graphs enable the analysis of a key property that we refer to as solvability. The identification of solvability guides better understanding and prediction of MARL interaction paradigms. To answer our research question, we utilize MAIDs to design the targeted intervention paradigm that intervenes on a single targeted agent, which reliably guides the entire multi-agent system towards a desired outcome. Formally, we frame this problem as selecting a desired Nash equilibrium among many possible equilibria harsanyi1988general , ensuring both primary task completion and satisfaction of an additional desired outcome.
The main contributions of this paper are summarized as follows: (1) We formalize the MARL learning process in the language of MAIDs. Based on MAIDs, we categorize the existing MARL interaction paradigms, and then propose a new one referred to as the targeted intervention paradigm. In this new interaction paradigm, only a single targeted agent is intervened on. (2) The solvability of existing MARL learning paradigms under various MARL interaction paradigms can be analyzed, with the help of visualizing the corresponding relevance graphs. (3) Since MAIDs can be viewed as causal diagrams, we draw on the principle of causal inference to implement the targeted intervention paradigm, which we refer to as the Pre-Strategy Intervention (PSI). It is provable that the PSI can reach a composite desired outcome consisting of the primary task goal and an additional desired outcome. (4) The PSI is implemented as a pre-policy module that can be integrated into generic MARL algorithms.
The proposed PSI is evaluated in Multi-Agent Particle Environment (MPE) and Hanabi lowe2017multi ; Bard_2020 . Furthermore, the solvability of different MARL learning paradigms concluded from relevance graphs are verified in experiments. Our code is publicly available as an open-source repository. [^1] https://github.com/iamlilAJ/Pre-Strategy-Intervention
2 Multi-Agent Influence Diagrams (MAIDs)
We now review multi-agent influence diagrams (MAIDs) (koller2003multi, ), which is an augmentation of the Bayesian network to describe multi-agent decision making to maximize their utility. In the rest of the paper, the terms variable and node will be used interchangeably. An MAID is usually described as a tuple $\mathcal{M}=(\mathcal{I},\mathcal{X},\mathcal{D},\mathcal{U},\mathcal{G},Pr)$, where $\mathcal{I}$ is a set of agents. $\mathcal{X}$ is a set of chance variables indicating decisions of nature. Each chance variable $X\in\mathcal{X}$ is associated with a set of parents $Pa(X)\subset\mathcal{X}\cup\mathcal{D}$. The $\mathcal{D}:=\bigcup_{i\in\mathcal{I}}\mathcal{D}_{i}$ is a set of all agents’ decision variables, where $\mathcal{D}_{i}$ is the set of agent $i$’s decision variables. For a decision variable $D\in\mathcal{D}_{i}$, $Pa(D)$ is the set of variables whose values are informed to agent $i$ when it selects a value of $D$. $\mathcal{U}:=\bigcup_{i\in\mathcal{I}}\mathcal{U}_{i}$ is a set of utility variables, where $\mathcal{U}_{i}$ is agent $i$’s utility variable as its utility function. Note that utility variables cannot be parents of other variables. A directed acyclic graph $\mathcal{G}$ with variables $\mathcal{V}=\mathcal{X}\cup\mathcal{D}\cup\mathcal{U}$ is formed. $Pr$ is a conditional probability distribution (CPD) defined over chance variables $X$ such as $Pr(X|Pa(X))$, and utility variables $U\in\mathcal{U}$ such as $Pr(U|\mathbf{pa})$, for each $\mathbf{pa}\in dom(Pa(U))$. Note that $Pr(U|Pa(U))$ is a Dirac function (i.e. $U$ is a deterministic function). In other words, for each instantiation $\mathbf{pa}\in dom(Pa(U))$, there is a value of $U$ that is assigned probability 1, and probability of other values is 0. To simplify the notation, $U(\mathbf{pa})$ denotes as the value of $U$ that has probability 1 when $Pa(U)=\mathbf{pa}$. The total utility that an agent $i$ obtained from an instantiation of $\mathcal{V}$ is the sum of the values of $\mathcal{U}_{i}$, i.e. $\sum_{U\in\mathcal{U}_{i}}U(\mathbf{pa})$ where $\mathbf{pa}\in dom(Pa(U))$.
Decision Rule and Strategy. An agent makes a decision at variable $D$ depending on its $Pa(D)$, determined by a decision rule $\delta:dom(Pa(D))\rightarrow\Delta(dom(D))$. $\Delta$ indicates the probability distribution space over a set. An assignment $\sigma$ of decision rules to each decision $D\in\mathcal{D}$ is called a strategy profile. A partial strategy profile $\sigma_{\mathcal{E}}$ is an assignment of decision rules to a subset of $\mathcal{D}$, as a restriction of $\sigma$ to $\mathcal{E}$, and $\sigma_{-\mathcal{E}}$ denotes the restriction of $\sigma$ to variables in $\mathcal{D}\backslash\mathcal{E}$. The assignment of $\sigma_{\mathcal{E}}$ to the MAID $\mathcal{M}$ induces a new MAID denoted by $\mathcal{M}[\sigma]$, and each $D\in\mathcal{E}$ would become a chance variable with the CPD $\sigma(D)$. When $\sigma$ is assigned to every decision variable in MAID, the induced MAID becomes a Bayesian network with no more decision variables. This Bayesian network defines a joint probability distribution $P_{\mathcal{M}[\sigma]}$ over all the variables in $\mathcal{M}$.
Expected Utility and Nash Equilibrium. Given a strategy profile assigned to each decision variable, with the resulting joint probability distribution $P_{\mathcal{M}[\sigma]}$ and suppose that $\mathcal{U}_{i}=\{U_{1},...,U_{m}\}$, we can write the expected utility for an agent $i$ as
$\mathbb{E}U_{i}(\sigma)=\sum_{(u_{1},...,u_{m})\in dom(\mathcal{U}_{i})}P_{\mathcal{M}[\sigma]}(u_{1},...,u_{m})\sum_{k=1}^{m}u_{k}.$ (1)
Given Eq. (1), the strategy $\sigma_{\mathcal{E}}^{*}$ is optimal for $\sigma$, for a subset $\mathcal{E}\subset\mathcal{D}_{i}$, if $\mathbb{E}U_{i}((\sigma_{-\mathcal{E}},\sigma_{\mathcal{E}}^{*}))\geq\mathbb{E}U_{i}((\sigma_{-\mathcal{E}},\sigma_{\mathcal{E}}^{\prime}))$, as shown in Definition 2.1. Furthermore, if for all agents $i\in\mathcal{I}$, $\sigma_{\mathcal{D}_{i}}$ is optimal for the strategy profile $\sigma$, then $\sigma$ is a Nash equilibrium (NE), as shown in Definition 2.2.
Definition 2.1 (Optimal Strategy koller2003multi ).
Let $\mathcal{E}$ be a subset of $\mathcal{D}_{i}$, and let $\sigma$ be a strategy profile. $\sigma_{\mathcal{E}}^{*}$ is optimal for the strategy profile $\sigma$ if, in the induced MAID $\mathcal{M}[\sigma_{-\mathcal{E}}]$, where the only remaining decisions are those in $\mathcal{E}$, the strategy $\sigma_{\mathcal{E}}^{*}$ is optimal, for all strategies $\sigma_{\mathcal{E}}^{\prime}$, such that
$\mathbb{E}U_{i}((\sigma_{-\mathcal{E}},\sigma_{\mathcal{E}}^{*}))\geq\mathbb{E}U_{i}((\sigma_{-\mathcal{E}},\sigma_{\mathcal{E}}^{\prime})).$
Definition 2.2 (Nash Equilibrium koller2003multi ).
A strategy profile $\sigma$ is a Nash equilibrium for a MAID $\mathcal{M}$ if for all agents $i\in\mathcal{I}$, $\sigma_{\mathcal{D}_{i}}$ is optimal for the strategy profile $\sigma$.
Remark 2.3.
[…]
Definition 2.4 (S-Reachability koller2003multi ).
A node $D^{\prime}$ is s-reachable from a node $D$ in a MAID $\mathcal{M}$ if there is some utility node $U\in\mathcal{U}_{D}$ such that if a new parent $\hat{D}^{\prime}$ were added to $D^{\prime}$, there would be an active path (see Example C.1 and Definition C.1 in Appendix C.2) in $\mathcal{M}$ from $\hat{D}^{\prime}$ to $U$ given $Pa(D)\cup\{D\}$, where a path is active in a MAID if it is active in the same graph, viewed as a Bayesian network (i.e. it is not blocked as per the standard d-separation rules). The relevance graph for a MAID $\mathcal{M}$ is a directed graph whose nodes are the decision nodes of $\mathcal{M}$, and which contains an edge $D^{\prime}\rightarrow D$ if and only if $D^{\prime}$ is s-reachable from $D$.
Relevance Graph. A relevance graph as shown in Definition 2.4 defines a directed graph induced from an MAID, describing the binary relation between two decision variables. If there exists an edge $D^{\prime}\rightarrow D$, it implies that the decision variable $D$ is strategically relies on another decision variable $D^{\prime}$. In other words, the decision rules for $D^{\prime}$ is required to evaluate the decision rules for $D$. If there exist both $D^{\prime}\rightarrow D$ and $D\rightarrow D^{\prime}$, then the relevance graph is cyclic. Furthermore, if $D$ and $D^{\prime}$ belong to two agents respectively, their payoffs depend on the decisions at both $D$ and $D^{\prime}$. Under this situation, the optimality of one agent’s decision rule is coupled with another agent’s decision rule.
Remark 2.5.
One way to figure out the above issue is making these two agents’ decision rules matched (koller2003multi, ), e.g. choosing both agents’ decision rules together, analogous to centralized training in multi-agent reinforcement learning (oliehoek2008optimal, ).
3 Multi-Agent Reinforcement Learning Interaction Paradigms
This paper focuses on analyzing interaction paradigms in MARL with guidance and proposes a novel approach for the targeted intervention paradigm. For a comprehensive understanding, we first formalize the two existing interaction paradigms: direct interaction and global intervention strouse2022collaborating ; mosqueira2023human ; wang2025m3hf , using multi-agent influence diagrams (MAIDs). Subsequently, we will elaborate on the details of our proposed targeted intervention paradigm.
[…]
$D_{h}$$D_{a}$
(a) Direct Interaction.
[…]
$D_{h}$$D_{a}$
(b) Global Intervention.
[…]
$D_{h}$$D_{a}$
(c) Targeted Intervention.
Figure 1: MAIDs (left part in each figure) and relevance graphs (right part) for one-shot simultaneous-move MARL interaction paradigms: (a) direct interaction, where agents interact directly; (b) global intervention, where a coordinator provides signals influencing all agents and their utilities; (c) targeted intervention, where a pre-strategy intervention applied on a targeted agent. In these diagrams, squares represent decision variables (specifically, red for the targeted agent, blue for other agent, yellow for a pre-decision); diamonds represent utility variables; and circles with $Z$ describe signals to guide agents for achieving their desired outcomes represented as utility variables. Relevance graphs depict strategic dependencies between these decision variables.
[…]
$D_{h}$$D_{a}$
(a) Direct Interaction.
[…]
$D_{h}$$D_{a}$
(b) Global Intervention.
[…]
$D_{h}$$D_{a}$
(c) Targeted Intervention.
Figure 2: MAIDs and their relevance graphs for one-shot sequential-move games. In addition to the variables introduced in Figure 1, the orange circles with II indicate the information sets.
3.1 Formalizing MARL Interaction Paradigms
By the definition of MAIDs in Section 2, we model MARL interaction paradigms based on agent variables ($\mathcal{I}$), decision variables ($\mathcal{D}$), chance variables ($\mathcal{X}$) representing states and information, and utility variables ($\mathcal{U}$) representing objectives. Utility variables capture the desired outcomes that agents pursue, whereas the special chance variables $\mathcal{Z}\subset\mathcal{X}$ describe signals (e.g. behaviours, communications, etc.) to guide agents for achieving their desired outcomes. The formal definition is delineated in Definition 3.1.
Definition 3.1.
A MARL interaction paradigm can be specified as a MAID $(\mathcal{I},\mathcal{X},\mathcal{D},\mathcal{U})$. $\mathcal{I}$ is the set of agents. $\mathcal{X}$ is the set of chance variables. $\mathcal{D}=\bigcup_{i\in\mathcal{I}}\mathcal{D}_{i}$ is the set of agent decision variables. $\mathcal{U}$ are utility variables representing objectives. The set of special chance variables ($\mathcal{Z}\subset\mathcal{X}$) describe signals to guide agents for achieving their desired outcomes.
Within the context of MARL, we define three interaction paradigms as follows:
(1) Direct Interaction. This paradigm represents the MARL where agents coordinate only via direct observation of the environment. There is no explicit, external coordination mechanism designed to steer agents towards their desired outcomes. Each agent must independently elicit a signal (a pattern of behaviours) based on local information to guide the other agent for achieving its desired outcome. For clarity of illustration, we focus on analyzing guidance to a single agent. We visualize this implicit process as a path from decision variables $D_{a}$ to the signal $Z$, and then to the utility nodes $U_{h}$, illustrated by MAID structures in Figure 1(a) and 2(a).
(2) Global Intervention. This paradigm features a centralized coordination providing simultaneous explicit guidance signals to all agents. In the MAID, as shown in Figure 1(b) and 2(b), this is demonstrated as information links from $Z$ to agents’ decision variables, and $Z$ usually influence their utility nodes (e.g., $U_{a}$ and $U_{h}$) directly, shaping their desired outcomes.
(3) Targeted Intervention. This paradigm involves applying an external guidance signal to a targeted agent, as shown in Figure 1(c) and 2(c). Specifically, $D_{pre}$ receives a guidance signal $Z$ and outputs filtered information. This intervention modifies the behaviours of the targeted agent, which in turn indirectly influences other agents. As a result, agents coordinate to achieve their desired outcomes provided by the guidance signal $Z$. While the concept of targeted intervention paradigm could potentially apply to multiple targeted agents with multiple desired outcomes, this paper specifically focuses on the case of a single targeted agent.
3.2 Interpreting MARL Interaction Paradigms
We now demonstrate how MAIDs interpret the three MARL interaction paradigms we mentioned above. Each MAID elucidates the structure of an interaction paradigm (through the role of variable $Z$) and its influence on decisions and utilities. Subsequently, we analyze how these structural differences translate into various strategic dependency patterns, as revealed by their relevance graphs.
Analyzing the relevance graph (Definition 2.4) corresponding to each MAID, reveals the key solvability property [^2] Seeking equilibria according to processing decision variables in s-reachability order (influencers first) koller2003multi . of MARL learning algorithms. As shown in Figure 1(a) and 2(a), the relevance graphs for the direct interaction paradigm are cyclic, implying computational or theoretical intricacies in finding solutions (Remark 2.5). In contrast, the relevance graphs for both the global intervention paradigm (Figures 1(b) and 2(b)) and targeted intervention paradigm (Figures 1(c) and 2(c)) are acyclic, suggesting they may be solvable by a broader class of MARL algorithms. It is possible in global intervention that agents learn independently based solely on the external central signal, in which case the relevance graph would have no edges at all, indicating no policy dependencies colas2022autotelicagentsintrinsicallymotivated . While both global intervention and targeted intervention paradigms offer acyclic structures, our proposed method relying on targeted intervention provides a distinct advantage: it achieves solvability and effectiveness only by intervention on a single targeted agent, offering a more practical approach in contrast to the global intervention.
3.3 Pre-Strategy Intervention: An Implementation of Targeted Intervention Paradigm
This section introduces the pre-strategy intervention—a principled design of the targeted intervention paradigm. Through the pre-strategy intervention, we aim to guide the multi-agent system towards a preferred Nash equilibrium (NE) that satisfies both the primary task goal and an additional desired outcome. Our approach first defines the pre-strategy intervention, then details how its influence on achieving this NE is quantified and optimized via learning a pre-policy to generate pre-strategies intervening on a single targeted agent.
3.3.1 Definition of Pre-Strategy Intervention
To regulate the intervened agent to maximize the desired outcome $U$, we propose to add a pre-decision variable $D_{pre}$ as a new parent to the decision variable $D_{h}$ of a selected agent $h\in\mathcal{I}$. In analogy to decision variables in MAIDs, we need to assign a strategy $\sigma_{pre}$ which we refer to as pre-strategy, and this operation is named pre-strategy intervention. This definition respects the convention of stochastic intervention in causal Bayesian networks (pearl2009causality, )[Chap. 4]. A pre-strategy is determined by a pre-policy denoted by $\delta_{pre}$, which processes the agent’s information and the guidance signal, as detailed in Definition 3.2.
Definition 3.2 (Pre-Strategy Intervention).
For a decision variable $D\in\mathcal{D}$ in a MAID, a pre-strategy intervention is an operation assigning a pre-strategy $\sigma_{pre}$ to a new parent $D_{pre}$ added to $D$, referred to as a pre-decision variable. The pre-strategy $\sigma_{pre}$ is determined by a pre-policy $\delta_{pre}:dom(Pa(D))\times\mathcal{Z}\rightarrow\Delta(dom(\sigma_{pre}))$, where $\mathcal{Z}$ is the space of guidance signals representing an additional desired outcome.
3.3.2 Causal Effect of Pre-Strategy Intervention
To design the pre-strategy intervention, we first define how its causal effect is measured. The composite desired outcome is captured by a total utility variable $U_{tot}:=U_{tot}^{sec}+U_{tot}^{task}$, comprising the primary task goal $U_{tot}^{task}$ and an additional desired outcome (a secondary goal) $U_{tot}^{sec}$. We formalize the causal effect of a pre-strategy intervention in Definition 3.3. The causal effect quantifies the total probabilities of $U_{tot}=u^{*}$, where $u^{*}$ is the optimal utility value of the desired composite outcome, under a pre-strategy intervention. In many cases, it is difficult to find a pre-strategy intervention that induces the optimal NE. Instead, we seek a pre-strategy that induces a probability distribution over a reduced set of NEs covering the optimal NE, denoted as $\boldsymbol{\hat{\sigma}}_{\mathcal{I}}$. Proposition 3.4 proves pre-strategy intervention maximizing this causal effect is guaranteed to exist.
Definition 3.3 (Causal Effect of Pre-Strategy Intervention).
A pre-strategy intervention is applied to a targeted agent’s strategy profiles, which influences the total utility $U_{tot}=u^{*}$ associated with a Nash equilibrium (NE) $\hat{\sigma}$ after the pre-strategy intervention. The set of all possible NEs is denoted by $\boldsymbol{\hat{\sigma}}$ (with no interventions). Given that $\boldsymbol{\hat{\sigma}}_{\mathcal{I}}$ denotes a reduced set of NEs induced by a pre-strategy intervention, covering the optimal NE (aligned to the additional desired outcome), the causal effect measuring this pre-strategy intervention is defined as follows:
$\begin{split}\Delta_{\text{CE}}^{\sigma_{pre}}(U_{tot}=u^{*})=\underbrace{\int_{\hat{\sigma}\in\boldsymbol{\hat{\sigma}}_{\mathcal{I}}}P_{\mathcal{M}[\hat{\sigma}]}(U_{tot}=u^{*})P_{\hat{\sigma}}(\hat{\sigma})\,d\hat{\sigma}}_{{P_{\mathcal{I}}}(U_{tot}=u^{*})}-\underbrace{\int_{\hat{\sigma}\in\boldsymbol{\hat{\sigma}}}P_{\mathcal{M}[\hat{\sigma}]}(U_{tot}=u^{*})P_{\hat{\sigma}}(\hat{\sigma})\,d\hat{\sigma}}_{P_{\mathcal{U}}(U_{tot}=u^{*})}.\end{split}$ (2)
[…]
Proposition 3.4.
Given a MAID $\mathcal{M}$, assume that the function $P_{\mathcal{I}}$ in Eq. (2), representing the probability of observing $U_{tot}=u^{*}$ under a pre-strategy intervention, is upper semicontinuous and defined on a compact domain $\operatorname{dom}(\sigma_{pre})\subseteq\mathbb{R}^{m}$. There exists at least a pre-strategy intervention on an agent that does not decrease the probability of $U_{tot}=u^{*}$. Furthermore, there exists a pre-strategy intervention that maximizes the causal effect.
To practically evaluate possible strategy profiles determined by the pre-strategies generated by a pre-policy, we reshape $P_{\mathcal{I}}$ as the following expression:
$\begin{split}P(U_{tot}=\ &u^{*}\mid\operatorname{do}(\sigma_{pre}))=\sum_{\hat{\sigma}\in\boldsymbol{\hat{\sigma}}}P_{\mathcal{M}[\hat{\sigma}]}(U_{tot}=u^{*})P_{\hat{\sigma}}(\hat{\sigma}\mid\operatorname{do}(\sigma_{pre})).\end{split}$ (3)
Note that $P_{\mathcal{U}}$ is a constant with respect to the intervention, so it can be ignored during optimization. In Eq. (3), $P_{\hat{\sigma}}(\hat{\sigma}\mid\operatorname{do}(\sigma_{pre}))$ is a probability distribution whose support includes all possible strategy profiles, which is feasible to implement and induce any possible $\boldsymbol{\hat{\sigma}}_{\mathcal{I}}$. Thus, we only need to maximize Eq. (3) in implementation.
4 MARL for Sequential Decision Making in MAIDs
We now show how the pre-strategy intervention introduced above is implemented in MARL for sequential decision making. We first show that each Markov game can be transformed into one multi-agent influence diagram (MAID). We model MARL using team reward Markov games littman1994markov . The main idea is establishing a connection that Markov games can be transformed into MAIDs, allowing our framework to apply. The formal definition of a Markov game and the detailed process of representing it as an MAID by matching their respective variables are provided in Appendix G. Building on this foundation, the subsequent subsections will extend the MARL interaction paradigms introduced before to Markov games (Section 4.1), discuss the solvability of MARL learning paradigms based on these MARL interaction paradigms (Section 4.2), and specify how the pre-strategy intervention in Eq. (3) is integrated into generic MARL learning algorithms (Section 4.3).
4.1 Extending MARL Interaction Paradigms to Markov Games
$D^{t}_{a}$$U^{t}$$D_{h}^{t}$$S^{t}$$U_{a}^{t}$$U_{h}^{t}$$Z^{t}$$S^{t+1}$
$D^{t}_{h}$$D^{t}_{a}$
(a) Direct Interaction.
$Z^{t}$$D^{t}_{h}$$D^{t}_{a}$$U^{t}$$S^{t}$$S^{t+1}$$U^{t}_{a}$$U^{t}_{h}$
$D^{t}_{h}$$D^{t}_{a}$
(b) Global Intervention.
$D^{t}_{h}$$D^{t}_{a}$$D^{t}_{\text{pre}}$$Z^{t}$$S^{t}$$U_{h}^{t}$$U^{t}$$U_{a}^{t}$$S^{t+1}$
$D^{t}_{h}$$D^{t}_{a}$
(c) Targeted Intervention.
Figure 3: MAIDs and their relevance graphs for dynamic simultaneous-move games. For conciseness, we only demonstrate one transition from timesteps $t$ to $t+1$. The variables used in one-shot decision making (see Figure  1) are decorated with timesteps. The semantics of these variables remain unchanged. The pink circles with $s_{t}$ and $s_{t+1}$ indicate states for timesteps $t$ and $t+1$. The green diamonds with $U^{t}$ indicate the total utility variables. The incoming edge from an individual utility variable ($U^{t}_{h}$ and $U^{t}_{a}$) to the total utility variable indicates that the agent’s individual utility contributes to the team reward.
$D^{t}_{a}$$I^{t}$$D^{t}_{h}$$S^{t}$$U^{t}_{a}$$U^{t}_{h}$$Z^{t}$$U^{t}$$S^{t+1}$
$D^{t}_{h}$$D^{t}_{a}$
(a) Direct Interaction.
$Z^{t}$$D^{t}_{h}$$D^{t}_{a}$$U^{t}$$S^{t}$$U^{t}_{a}$$U^{t}_{h}$$I^{t}$$S^{t+1}$
$D^{t}_{h}$$D^{t}_{a}$
(b) Global Intervention.
$S^{t}$$D^{t}_{h}$$D^{t}_{a}$$I^{t}$$D^{t}_{\text{pre}}$$Z^{t}$$U_{h}^{t}$$U^{t}$$S^{t+1}$$U_{a}^{t}$
$D^{t}_{h}$$D^{t}_{a}$
(c) Targeted Intervention.
Figure 4: MAIDs and their relevance graphs for dynamic sequential-move games. For conciseness, we only demonstrate one transition from timesteps $t$ to $t+1$. In addition to the variables introduced in Figure 3, the orange circles with $I^{t}$ indicate the information sets.
We now demonstrate how to extend the three MARL interaction paradigms (see Definition 3.1) to Markov games through the lens of MAIDs. Each interaction paradigm under sequential decision making in dynamic environments with simultaneous or sequential moves can be extended from one-shot decision making introduced in Section 3.2, by considering timesteps and states. With representation of the total utility variable, we can derive the MAID graphical representation of these three interaction paradigms in Markov games, which are illustrated in Figures 3 and 4. Recall that the variable $\boldsymbol{Z^{t}}$ has different semantic meanings across different MARL interaction paradigms, similar to the variable $\boldsymbol{Z}$ in Definition 3.1.
It can be observed that the widespread global intervention paradigm implemented for MARL algorithms (e.g., applying a centralized coordinator wang2020roma ; na2024lagma ; liu2023lazy ) can be justifiably interpreted with a visualization. More importantly, it opens a new door to design novel MARL interaction paradigms to solve a broad class of problems with the help of MAIDs, such as the targeted intervention paradigm proposed in this paper to solve the team reward Markov games.
4.2 Solvability of MARL Learning Paradigms under Various Interaction Paradigms
Independent Learning. Independent learning (IL) algorithms  [^3] Note that the IL algorithms we mention here exclude those with additional mechanisms designed to improve convergence, such as the optimistic mechanism (lauer2000algorithm, ). are the direct application of single-agent reinforcement learning algorithms, such as temporal-difference (TD) learning (sutton2018reinforcement, )[Chap. 6] and the actor-critic algorithms (konda1999actor, ). If a relevance graph is cyclic, it is not guaranteed to reach an NE by solely determining each agent’s decision variables with a generalized backward induction algorithm in analogy to independent learning, as discussed in Section 2.4. This aligns with the direct interaction paradigm (Figure 3(a) and 4(a)), reflecting the non-stationarity dilemma (hernandez2019survey, ). Equipping IL algorithms with the global intervention (Figures 3(b) and 4(b)) and the targeted intervention (Figures 3(c) and 4(c)) paradigms can generate acyclic relevance graphs. Such acyclic relevance graphs can guarantee the tampered IL algorithms reaching an NE with theoretical guarantees.
Centralized Training. As discussed in Remark 2.5, the only way to address cyclic relevance graphs is enabling two agents’ decision making to be matched. One solution is transforming a cyclic relevance graph to a component graph, where each maximal strongly connected component (SCC) is regarded as a supernode de2023graph . More specifically, a maximal SCC consists of all the decision variables forming a cyclic relevance graph at each timestep. Koller and Milch koller2003multi showed that solving the acyclic component graph [^4] A component graph is always acyclic (cormen2022introduction, ). by the generalized backward induction algorithm can reach an NE. This is in coincidence associated with the centralized training employed to reach the maximum cumulative team rewards in Markov games (oliehoek2008optimal, ; littman2001value, ). In summary, centralized training can be seen as an algorithmic approach to address the issues of pure IL algorithms existing in the direct interaction paradigm, rather than switching to other interaction paradigms.
4.3 Pre-Policy Learning in MARL
As we have established how MARL interaction paradigms can be represented in MAIDs and subsequently how it can fit in the team-reward Markov games, the techniques developed in Section 3.3.2 can guide us to realizing the pre-strategy intervention in the team-reward Markov games. Following Definition 3.2, the requirement is to learn a pre-policy $\delta_{pre}:dom(Pa(D))\times\mathcal{Z}\rightarrow\Delta(dom(\sigma_{pre}))$ to generate the pre-strategy $\sigma_{pre}$ for the possible optimal NEs. This can be realized by maximizing the causal effect expressed in Eq. (3). Our proposed algorithm, which can be integrated with generic MARL algorithms, is referred to as Pre-Strategy Intervention (PSI).
Practically, Eq. (3) is implemented by Monte Carlo sampling, the RHS of which can be reformulated as a posterior distribution over a Nash equilibrium $\sigma$, as follows:
$P(\sigma\mid U_{tot}=u^{*},\operatorname{do}(\sigma_{pre}))\propto P_{\mathcal{M}[\sigma]}(U_{tot}=u^{*}\mid\sigma)P_{\sigma}(\sigma\mid\operatorname{do}(\sigma_{pre})).$ (4)
Maximizing the posterior distribution $P(\sigma\ |\ U_{tot}=u^{*},\operatorname{do}(\sigma_{pre}))$ in Eq. (4) reveals a fact that observing the total utility variable $U_{tot}$ (rewards) and a pre-strategy intervention $\operatorname{do}(\sigma_{pre})$ are the key conditions for seeking the optimal joint agent strategy profile denoted by $\sigma$. This more intuitively illustrates the maximization of the causal effect of the pre-strategy intervention in Eq. (3), which provides a clear probabilistic objective that can be optimized by generic MARL algorithms.
5 Experiments
5.1 Experimental Setups
We evaluate our proposed Pre-Strategy Intervention (PSI) in two environments: the Multi-Agent Particle Environment (MPE) lowe2017multi , a simultaneous-move cooperative navigation game, and the Hanabi card game Bard_2020 , which challenges coordination within a multi-agent system under partial observability. Hanabi is known for having multiple, distinct equilibria, making it an ideal testbed for analyzing convergence to a specific Nash equilibrium. In both settings, the intervention is applied to a single targeted agent to assess its effectiveness in achieving the additional desired outcome and improving the task completion. For the experiments presented in this paper, we select a fixed agent for intervention in each environment, which allows for a controlled analysis of the PSI. Full details of these environments, including specific experimental configurations and how the additional desired outcomes are defined and implemented, are provided in Appendix H. In all environments, intrinsic rewards are used for measuring reachability of the additional desired outcomes, while extrinsic rewards are used for measuring the task completion. Results from 5 random seeds are reported as means with $95\%$ confidence intervals.
Implementation of Our Method. Our PSI is implemented by designing a pre-policy module (a GRU or MLP, matching agent’s backbones) to realize the targeted intervention paradigm, which takes a concatenated input of environmental observations and a measure representing an additional desired outcome (formulated as an intrinsic reward). [^5] This implementation maintains solvability of the PSI (see Appendix F.2). It outputs an embedding, which is then fed to Q-value functions for value-based MARL methods or critics for policy-based methods (see Appendix H for details).
Baselines. All baselines are implemented with the same architecture and training setups as our method. We evaluate our method against several base MARL algorithms referred to as Base MARL (direct interaction paradigms): for MPE, these include DQN (mnih2015human, ), VDN (sunehag2017value, ) and QMIX (rashid2020monotonic, ); for Hanabi, we use IPPO (schulman2017proximal, ), MAPPO (yu2022surprising, ) and PQN (gallici2024simplifying, ) (with PQN-IQL for independent learning and PQN-VDN for centralized training variants). An ablation study that removes the pre-policy module while retaining the intrinsic reward maximization examines the effectiveness of the targeted intervention paradigm without the confounding effect of the pre-policy module. This ablated variant is referred to as Intrinsic Reward that still satisfies the targeted interaction paradigm, with removing pre-decision variables $D_{pre}$ in Figures 3 and 4 and connecting the $Z^{t}$ interpreted as the intrinsic rewards to maximize directly to the decision variables $D$. To emphasize the strength of our targeted intervention paradigm, we also compare against the existing approaches implemented in the global intervention paradigm, modifying LIIR du2019liir or LAIES liu2023lazy as representative policy-based or value-based methods. This baseline is referred to as Global Intervention.
5.2 Results and Analysis
(a) MPE: Main results (extrinsic left, intrinsic right).
(b) MPE: Targeted vs. Global
(c) Hanabi: Main results (extrinsic left, intrinsic right).
(d) Hanabi: Targeted vs. Global
Figure 5: In Subfigures a and c, our method PSI is compared against Intrinsic Reward and ase MARL; these plots show both extrinsic (task completion) and intrinsic (additional desired outcome) returns. Our targeted intervention paradigm (PSI) is compared against the methods of global intervention paradigm in Subfigures b and d with extrinsic returns.
Additional Goals Improving Task Performance. As Figures 5(a) and 5(c) show, both our PSI and the Intrinsic Reward baseline outperform base MARL algorithms. These results confirm that incorporating guidance from an additional desired outcome enhances task completion.
Verification of Relevance Graph Analysis. Our MAID and relevance graph analysis predicts that PSI improves the solvability of MARL learning paradigms (Section 4.2), enabling the pure independent learning more solvable. Experiments verify this across different settings. In simultaneous-move MPE (analysis in Figure 3(c) and results in Figure 5(a)), Independent Q-Learning (IQL) equipped with our PSI achieves task completion comparable to VDN, a centralized training method. This is a notable improvement for IQL over its baseline performance (without our PSI). Similarly, in sequential-move Hanabi (analysis in Figure 4(c) and results in Figure 5(c)), independent learning algorithms augmented with the PSI achieve performance comparable to, or even better than centralized training.
Targeted Intervention Outperforming Global Intervention. Our targeted intervention paradigm implemented in PSI consistently outperforms the global intervention paradigm across various MARL backbones, as shown in Figures 5(b) (MPE) and 5(d) (Hanabi). While both paradigms exhibit acyclic relevance graphs suggesting that the corresponding MARL algorithms are solvable in theory, [^6] This can be observed from the MAID analyses in Figures 3(c) and 4(c) for the targeted intervention paradigm, and Figures 3(b) and 4(b) for the global intervention paradigm. the global intervention paradigm’s empirical performance often suffers. We attribute this to the inherent practical difficulty of designing (learning) an effective coordination mechanism to globally assign beneficial, non-conflicting subgoals to multiple agents simultaneously. This can subsequently hinder their learning of optimal independent policies.
Ablation: Our Method vs. Intrinsic Reward. To assess our pre-policy module’s contribution, we conduct an ablation study to compare our PSI against the Intrinsic Reward baseline, both of which belong to the targeted intervention paradigm. Experimental results presented for MPE in Figure 5(a) and for Hanabi in Figure 5(c) reveal a key distinction: while both our method and the Intrinsic Reward baseline effectively lead to the attainment of the additional desired outcome (reflected in the high intrinsic return), our method with the pre-policy module achieves notably superior performance in primary task completion (reflected in the higher extrinsic return). This highlights that the pre-policy module is instrumental for retaining the targeted agent’s adherence to an additional desired outcome while not forgetting the task goal.
Analysis of Nash Equilibrium Convergence in Hanabi. The Hanabi environment, with its multiple distinct equilibria Bard_2020 , serves as an ideal testbed for analyzing convergence to a desired Nash equilibrium. We treat a specific human convention as a desired, high-performing Nash equilibrium (NE) and design the intrinsic reward to directly measure the agents’ compliance to it. As shown in Figure 5(c), the high and stable intrinsic return achieved by our targeted intervention paradigm provides a strong evidence that the agents successfully converge to the desired NE. In contrast, the low intrinsic return of the baseline methods suggest they fail to establish a coordinated convention and are likely stuck in one of Hanabi’s many inferior equilibria.
Appendix I details further exploration of our method’s capabilities and robustness. These include: (1) evaluation using varied additional desired outcomes in MPE and Hanabi; and (2) performance assessments under noisy observation conditions.
6 Conclusion
Summary. We incorporate multi-agent influence diagrams (MAIDs) as a principled framework into designing and analyzing the targeted intervention paradigm in MARL. Our Pre-Strategy Intervention (PSI) as an implementation of the targeted intervention paradigm, applied to a single targeted agent, maximizes the causal effect on achieving a composite desired outcome that integrates both the primary task goal and an additional desired outcome. Furthermore, the relevance graphs of MAIDs offer theoretical insights into the solvability of MARL learning paradigms under various interaction paradigms. Experiments verify the effectiveness of the PSI.
Limitation. Our principle currently presumes that the underlying structure of MARL interaction paradigms in MAIDs is complete or can be precisely modelled. This structural knowledge is a prerequisite for effective influence propagation via targeted intervention and can be challenging to define a priori in realistic environments or complex systems. Additionally, our analysis primarily focuses on a single targeted intervention, which may limit its applicability to real-world settings.
Future Work. Key future directions of our work include but are not limited to:
  1. 1.
    Learning MAID structures from data (e.g., via causal discovery kenton2023discovering ) to reduce reliance on pre-specified models. Further, the MAID can be seen as a foundation to describe the world model.
  2. 2.
    Coordinating a multi-agent system under the targeted intervention paradigm on multiple agents.
  3. 3.
    Designing the optimal criteria for selecting the number of targeted agents and appointing proper agents with various types to complete different tasks in various scenarios. The information-theoretic measure of empowerment jaques2019social ; klyubin2005empowerment offers a potential quantitative basis for such criteria.
  4. 4.
    Integrating advanced reasoning modules (e.g. large language models and the broader generative AI) to enhance the capability of the PSI on the targeted agent in online learning and online adaptation to unknown agent teammates and unseen situations (wang2024open, ).
  5. 5.
    The principle proposed in this paper can be used to design new MARL interaction paradigms.
Acknowledgment
Jianhong Wang is supported by the Engineering and Physical Sciences Research Council (EPSRC) [Grant Ref: EP/Y028732/1]. Samuel Kaski is supported by UKRI Turing AI World-Leading Researcher Fellowship, EP/W002973/1.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We raise our research question and solve it then evaluate.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We have well discussed the limitations in Section 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The omitted proof of main paper is in the Appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide implementation details in the Appendices.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The code for reproducing the main experimental results is published on: https://github.com/iamlilAJ/Pre-Strategy-Intervention.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide details of the experimental settings in Appendices.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Results are shown with means and $95\%$ confidence intervals.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We disclose the computational resources needed in Appendices.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our research complies with the NeurIPS Code of Ethics. The paper presents algorithmic work in simulated environments without human subjects or potential for immediate harmful applications.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Our work focuses on foundational research in multi-agent reinforcement learning and introduces a new theoretical framework for agent coordination. The proposed Pre-Strategy Intervention method remains at a theoretical and algorithmic level, with experiments conducted only in simulated environments (MPE and Hanabi).
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our paper does not release models or datasets with high risk for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have cited the project we used.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: All code is extensively commented.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our research does not involve any crowdsourcing or human subjects. All experiments were conducted in simulated environments.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our research does not involve human subjects or participants, as all experiments were conducted in simulated reinforcement learning environments.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLMs were only used for grammar checking and text formatting.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Notation
Table 1: Summary of notations used in this paper.
Notation Description
MAID Core Components
$\mathcal{M}$ Multi-Agent Influence Diagram (MAID).
$\mathcal{I}$ Set of agents in the MAID.
$\mathcal{X}$ Set of chance variables (representing states, information, etc.).
$\mathcal{D},\mathcal{D}_{i}$ Set of decision variables for all agents, and for agent $i$.
$\mathcal{U},\mathcal{U}_{i}$ Set of utility variables (objectives) for all agents, and for agent $i$.
$\operatorname{Pa}(D)$ Parent set of a decision variable $D$.
$\mathcal{G}$ Directed acyclic graph (DAG) of the MAID.
PrPr Probability distribution.
$\sigma$ A strategy profile (assignment of decision rules to all agents).
$\hat{\sigma}$ An arbitrary Nash equilibrium (NE).
$\boldsymbol{\hat{\sigma}}$ The set of all possible NEs.
Pre-Strategy Intervention
$\mathcal{Z}\subset\mathcal{X}$ The space of guidance signals representing all possible additional desired outcomes.
$Z\in\mathcal{Z}$ A specific guidance signal representing an additional desired outcome.
$D_{pre}$ Pre-strategy decision variable.
$\sigma_{pre}$ A pre-strategy assigned to the pre-decision variable $D_{pre}$.
$\delta_{pre}$ A pre-policy, which is a function that determines a pre-strategy $\sigma_{pre}$.
$\boldsymbol{\hat{\sigma}}_{\mathcal{I}}$ The reduced set of NEs induced by an intervention.
Utility, Probability Distribution and Causal Effect
$U_{tot}$ Total utility, representing the overall objective.
$U_{tot}^{task},U_{tot}^{sec}$ Utility of the primary task goal and the secondary goal, respectively.
$\mathbb{E}_{U_{i}}(\sigma)$ Expected utility for agent $i$ under strategy profile $\sigma$.
$P_{\mathcal{M}[\sigma]}(U=u)$ Likelihood of outcome $U=u$ under strategy profile $\sigma$.
$P_{\sigma}(\sigma)$ Probability distribution over strategy profiles.
$P(\cdot\mid\operatorname{do}(\sigma_{pre}))$ Probability distribution after applying the causal intervention $\operatorname{do}(\sigma_{pre})$.
$\Delta_{\text{CE}}^{\sigma_{pre}}(U=u)$ Causal effect of a pre-strategy intervention on an outcome.
MARL
$\mathcal{S}$ Set of states in a Markov game.
$\mathcal{A}$ Joint actions in a Markov game.
$T$ Transition function in a Markov game.
$R$ Team reward function in a Markov game.
$\pi$ Joint policy of agents in a Markov game.
Appendix B Related Work
B.1 Multi-agent Team with Human Feedback
Our work connects to the growing field of leveraging human guidance within multi-agent systems [44]. A prominent related area is reinforcement learning from human feedback (RLHF), where human input, such as preferences or demonstrations, guides agent policy optimization [45, 46], often proving essential for tasks with complex or subjective goals for which explicit reward design is difficult [46, 47, 48]. RLHF is a specific instance within the broader scope of human-instructed multi-agent systems, which utilize diverse inputs like direct commands [49], demonstrations [50, 51], policy shaping [47], or language instructions [52] to steer agent behaviors. Incorporating such human expertise can significantly enhance system performance [45, 53, 54, 55], interpretability [11, 52, 56, 57], and alignment [46, 58]. These methods typically treat human feedback as a data source to directly train an agent’s policy, often in a model-free manner.
In contrast, our work introduces a structural approach grounded in MARL interaction paradigms. We use multi-agent influence diagrams (MAIDs) to formally model the strategic dependencies within a MARL learning process. Based on this model, our targeted intervention paradigm is designed to provably maximize the causal effect on a desired outcome, which can be provided by human feedback. The core advantage of our framework is therefore its ability to provide a principled, analytical method for designing an intervention on a multi-agent system and predicting its system-level impact. This is a significant advantage in scenarios where the effects of global, heuristic feedback are hard to foresee.
B.2 Goal-Conditioned MARL
Our work relates to goal-conditioned reinforcement learning (GCRL), where agents learn policies conditioned on achieving diverse goals, such as reaching specific states [59, 60], matching images [61], or following language instructions [62]. While concepts similar to subgoals like roles [27], skills [63], or subtasks [64] are used in MARL for task decomposition or coordination. The foundational approach in this area is the option framework in hierarchical reinforcement learning (HRL) [65], where “options” (temporally extended sub-policies) are used to achieve subgoals.
Our “additional desired outcome” can serve a similar function to a subgoal, though our framework interprets this concept more broadly. A traditional subgoal often refers to a specific intermediate state required for task completion. However, our composite desired outcome can also encompass beneficial patterns of behaviour that, while not directly related to the task completion, significantly improve the MARL learning performance and coordination. For example, in the game Hanabi, a team can achieve a score without adhering to a specific convention. However, adopting a convention (an additional desired outcome) makes teammates’ actions more predictable, thereby enhancing coordination and improving overall performance.
This broader focus on guiding patterns of behaviour, rather than just reaching explicit, predefined subgoals, motivates our distinct approach. Our approach differs from these methods in two principled ways. First, our Pre-Strategy Intervention provides a continuous guidance signal that influences an agent’s primitive actions at each timestep, rather than directly invoking a temporally abstract sub-policy (e.g. option) to make decision. Furthermore, our approach is designed to be particularly advantageous in scenarios where providing global guidance to all agents is impractical. Our core distinction lies in using a principled framework to analyze and apply a targeted intervention to a single agent.
B.3 Intrinsic Reward Method
A common approach to realize guidance in MARL is through maximizing intrinsic rewards during learning. These internally generated signals can encourage diverse behaviours motivated by curiosity [66, 67], novelty [68, 69], social influence [7], empowerment [9], learned reward functions [8, 70] or utilizing prior knowledge from large language models [23, 52].
While our Pre-Strategy Intervention is implemented via an intrinsic reward, our core distinction lies in its purpose and scope. Unlike methods such as those in  [8, 23], which applied intrinsic rewards to all agents to encourage emergent coordination, our Pre-Strategy Intervention strategically directs the signal only to a single targeted agent. Furthermore, standard intrinsic rewards are generally designed to solve local challenges like sparse rewards via using heuristic metrics like novelty. In contrast, our guidance signal is explicitly designed to solve a system-level problem: reaching a desired outcome for the entire multi-agent system.
To achieve this, we leverage Multi-Agent Influence Diagrams (MAIDs) and their corresponding relevance graphs to formally analyze the strategic dependencies between agents. This causal analysis is crucial as it allows us to move beyond simple heuristics. We then strategically design a guidance signal, which is implemented as an intrinsic reward for a single targeted agent. The specific objective of this signal is to maximize its causal effect on the composite desired outcome.
Therefore, our contribution is not merely the use of an intrinsic reward, but rather the principled framework that uses causal analysis to inform the design of such a signal.
B.4 Environment and Mechanism Design
Our work relates to concepts from environment and mechanism design. Environment design typically involves structuring or modifying environmental configurations to guide agent behaviors towards a targeted outcome [71, 72, 73]. In contrast, our approach does not directly reconfigure the static environment. Rather, it focuses on intervening on a targeted agent. Consequently, the behavior of the intervened agent changes, altering the effective environment dynamics experienced by other agents. This allows us to shape emergent team behavior by modifying agent interaction paradigms rather than fixed environmental rules.
Our work can also be viewed through the lens of mechanism design, which focuses on designing the “rules of games” such that a desired outcome is attained [74, 75]. Our proposed targeted intervention paradigm can be categorized into this context as a mechanism that is applied to specific agents designed to steer the MARL learning process towards equilibria that satisfy both the primary task goal and a specified additional desired outcome.
B.5 Probabilistic Graphical Models for Multi-Agent Games
Graphical models based on Bayesian networks were introduced to represent dependencies, relevance, and relationships among variables in multi-agent games by Koller and Milch [16]. This formulation has since inspired advancements in both algorithmic innovations for solving games [76, 77] and empirical analysis [78]. More recently, Hammond et al. [79] incorporated causal graphical models into multi-agent reasoning, extending multi-agent influence diagrams (MAIDs) to the causal domain. Their work introduced the concept of pre-policy intervention, allowing for queries such as “What if some agents commit to a policy before others make their decisions?” This framework has become a foundational tool in subsequent research on theory of mind [80, 81], reinforcement learning [82], and causal modeling of agents [41]. To maintain consistency with the MAID framework, we rename “pre-policy intervention” to “pre-strategy intervention” in this paper.
Our work builds directly on the causal framework proposed by Hammond et al. [79]. However, whereas their work introduced pre-policy intervention as a general tool for causal reasoning in game theory, our work is the first to adapt and apply this concept to solve a core challenge in MARL: achieving coordination by guiding a multi-agent system towards a desired outcome. We leverage MAIDs not just for analysis, but as a practical design tool to construct an MARL interaction paradigm. We design Pre-Strategy Intervention as the core mechanism within the targeted intervention paradigm, using it to guide the MARL learning process towards a desired, high-performing outcome.
Appendix C Extra Background
C.1 An Example of MAIDs
Poison TreeTree DoctorBuild PatioTree SickCostEffortTree DeadTreeView
Figure 6: A MAID for the Tree Killer example; Alice’s decision and utility variables are in red, and Bob’s are in blue. Decision nodes are rectangular, chance nodes are squircular, and utility nodes are hexagonal.
We exemplify MAIDs through a two-agent scenario from [16].
Example.
Alice is considering building a patio behind her house, which would be more valuable if she could have a clear view of the ocean. However, a tree in the yard of her neighbor Bob blocks her view. Alice, being somewhat unscrupulous, contemplates poisoning Bob’s tree, which would cost her some effort but might cause the tree to become sick. Bob is unaware of Alice’s actions, but can observe if the tree starts to deteriorate, and he has the option of hiring a tree doctor (at a cost). The attention of the tree doctor reduces the chance that the tree will die during winter. Meanwhile, Alice must decide whether to build her patio before the weather turns cold. At the time of her decision, Alice knows whether a tree doctor has been hired but cannot directly observe the tree’s health. An MAID for this scenario is shown in Figure 6.
In this example, Poison Tree and Tree Doctor are s-reachable to Build Patio, as Alice must take into account whether she has already poisoned the tree and whether Bob has called a tree doctor before deciding whether to build the patio.
C.2 Additional Definitions
Definition C.1 (Active Path [83]).
Let $\mathcal{G}$ be a Bayesian Network (BN) structure, and let $X_{1}-X_{2}-\dots-X_{n}$ represent an undirected path in $\mathcal{G}$. Let $E$ be a subset of nodes in $\mathcal{G}$ (the evidence set). The path $X_{1}-\dots-X_{n}$ is active given evidence $E$ if the following conditions hold:
  1. 1.
    Whenever a collider is on the path, i.e., a structure $X_{i-1}\to X_{i}\leftarrow X_{i+1}$, then either $X_{i}$ or one of its descendants is in $E$;
  2. 2.
    No other nodes along the path is in $E$.
Appendix D Illustrative Example: Pre-strategy Intervention in a Logistics Game
Background
Two logistics companies, Company A and Company B, share a warehouse and use robots to manage inventory. Each company has two options:
  • Optimize space usage: Focus on efficient organization.
  • Prioritize speed: Focus on moving items quickly.
Both companies’ choices affect each other’s performance, and they aim to achieve the best outcome for their operations.
Utility Table
[…]
D.1 Pre-Strategy Intervention
Our Pre-Strategy Intervention is applied to Company A (the targeted agent) to guide the system towards the more desired (9,9) equilibrium. The pre-policy module learns to provide a guidance signal to Company A that effectively incentivizes it to choose “Optimize Space Usage.” For instance, this guidance might be implemented as an additional input signal that favorably alters Company A’s perceived utility or Q-values for choosing “Optimize Space Usage,” making it the optimal choice for Company A given the intervention.
By influencing Company A to select “Optimize Space Usage,” and assuming Company B best responds to this action (or also converges to “Optimize Space Usage” due to the system dynamics and its own objectives), the system is steered towards the (9,9) outcome. This targeted intervention on Company A thus fosters a cooperative outcome and prevents convergence to the less efficient (5,5) equilibrium, demonstrating how Pre-Strategy Intervention can be used for equilibrium selection in a MARL setting.
Appendix E Theoretical Proof of Proposition 3.4
The proposition relies on the assumption that the probability function $P_{\mathcal{I}}$ is upper semicontinuous. This is a reasonable assumption because the best-response function in a game is not always continuous. An intuitive example to support this assumption is the game paper, rock, scissors, where the best response is conducting each action uniformly. If we consider a pre-strategy that shifts one player towards slightly less likely playing rock, then the probability of the opponent playing paper would experience a “jump” to 0, which can be seen as a discontinuity in the function. A detailed example solution of pre-strategy intervention can be found in Appendix D.
Proof.
Proposition 3.4 claims that an intervention exists: (a) it does not decrease the probability of the desired outcome and (b) a maximum-effect intervention exists.
(a) A trivial case exists where a pre-strategy that equals the marginal conditional probability of $U=u$ can be achieved by doing empty intervention.
(b) To prove that there exists a pre-strategy maximizing the causal effect, we observe that the second term on the right-hand side of Equation (2) is constant. Therefore, maximizing the first term is equivalent to maximizing the causal effect.
To analyze how an agent’s strategy influences the final outcome, we can decompose the probability of achieving the desired outcome $U_{tot}=u^{*}$. Using the law of total probability and the Markov property of the Bayesian network [16], we can marginalize out the decision variable $D$ and its parents $Pa(D)$. This decomposition explicitly isolates the agent’s decision rule, $P_{\mathcal{M}[\sigma]}(d\mid\mathbf{pa}_{D})$, which is the component that our intervention targets:
$P_{\mathcal{M}[\sigma]}(U_{tot}=u^{*})=$ $\int_{\mathbf{pa}_{D}\in\operatorname{dom}(Pa(D))}P_{\mathcal{M}[\sigma]}(\mathbf{pa}_{D})\,d\mathbf{pa}_{D}$
$\times\int_{d\in\operatorname{dom}(D)}P_{\mathcal{M}[\sigma]}(d\mid\mathbf{pa}_{D})\,dd$
$\times P_{\mathcal{M}[\sigma]}(U_{tot}=u^{*}\mid d,\mathbf{pa}_{D})$ (5)
The function $f(\sigma_{pre})$, representing the expected probability of $U=u$ under the pre-strategy, is defined as:
$f(\sigma_{pre}):=P_{\mathcal{I}}(U_{tot}=u^{*})=\int_{\hat{\sigma}\in\boldsymbol{\hat{\sigma}}_{\mathcal{I}}}P_{\mathcal{M}[\hat{\sigma}]}(U_{tot}=u^{*})P_{\hat{\sigma}}(\hat{\sigma})\,d\hat{\sigma}$
Assuming $f$ is an upper semicontinuous function defined on a compact domain $\operatorname{dom}(\sigma_{pre})\subseteq\mathbb{R}^{m}$, we aim to demonstrate that $f$ has a maximum on this domain. This follows from the Extreme Value Theorem. We may replace the notation $\sigma_{pre}$ with $\sigma$ in the following steps for simplicity, with a slight abuse of notation.
Boundedness Above: Suppose, for contradiction, that $f$ is unbounded above. For each $k\in\mathbb{N}$, there exists $\sigma_{k}\in\operatorname{dom}(\sigma)$ such that $f(\sigma_{k})>k$. Since $\operatorname{dom}(\sigma)$ is compact, the sequence $\{\sigma_{k}\}$ contains a convergent subsequence $\{\sigma_{k_{l}}\}$ converging to some $\sigma_{0}\in\operatorname{dom}(\sigma)$.
The property of upper semicontinuity implies $\limsup_{l\to\infty}f(\sigma_{k_{l}})\leq f(\sigma_{0})$, which contradicts the assumption because it suggests $\limsup_{l\to\infty}f(\sigma_{k_{l}})=\infty$. This shows f is bounded above. Then we can define:
$\gamma=\sup\{f(\sigma):\sigma\in\operatorname{dom}(\sigma)\}$
Since the set $\{f(\sigma):\sigma\in\operatorname{dom}(\sigma)\}$ is nonempty and bounded above, $\gamma\in\mathbb{R}$.
Existence of Maximum: Let $\left\{\sigma_{k}\right\}$ be a sequence in $\operatorname{dom}(\sigma$ such that $\left\{f\left(\sigma_{k}\right)\right\}$ converges to $\gamma$. By the compactness of the domain, the sequence $\left\{\sigma_{k}\right\}$ has a convergent subsequence $\left\{\sigma_{k_{\ell}}\right\}$ that converges to some $\bar{\sigma}\in\operatorname{dom}(\sigma)$. Then
$\gamma=\lim_{\ell\rightarrow\infty}f\left(\sigma_{k_{\ell}}\right)=\limsup_{\ell\rightarrow\infty}f\left(\sigma_{k_{\ell}}\right)\leq f(\bar{\sigma})\leq\gamma$
Conclusion: The equality $\gamma=f(\bar{\sigma})$ establishes that $\gamma$ is the maximum value of $f$ on $\operatorname{dom}(\sigma)$, and thus $f(\sigma)\leq f(\bar{\sigma})$ for all $\sigma$ in the domain $\operatorname{dom}(\sigma)$.
Appendix F Multi-Agent Influence Diagrams and Relevance Graphs in Implementation
$D_{h}$$U_{h}$$D_{\text{pre}}$$Z_{h}$$D_{a}$$U_{a}$
$D_{h}$$D_{a}$
(a) Simultaneous Moves.
[…]
$D_{h}$$D_{a}$
(b) Sequential Moves.
Figure 7: MAIDs and relevance graphs for MARL interaction paradigms. The figures illustrate the structure for (a) Simultaneous moves and (b) Sequential moves. The variable “ZZ” here indicates an intrinsic reward to implement an additional desired outcome.
$D^{t}_{h}$$D^{t}_{a}$$D^{t}_{\text{pre}}$$Z^{t}$$S^{t}$$U_{h}^{t}$$U^{t}$$U_{a}^{t}$$S^{t+1}$
$D^{t}_{h}$$D^{t}_{a}$
(a) Simultaneous Moves.
$S^{t}$$D^{t}_{h}$$D^{t}_{a}$$I^{t}$$D^{t}_{\text{pre}}$$Z^{t}$$U_{h}^{t}$$U^{t}$$S^{t+1}$$U_{a}^{t}$
$D^{t}_{h}$$D^{t}_{a}$
(b) Sequential Moves.
Figure 8: MAIDs for pre-strategy intervention approach implemented in experiments, in sequential-move sequential decision makings, their corresponding Markov games and the relevance graphs for the Markov games. The variable $Z^{t}$” here indicates an intrinsic reward to implement an additional desired outcome.
In this section, we provide a more detailed illustration of the multi-agent influence diagram (MAID) structure that represents the practical implementation of our proposed Pre-Strategy Intervention. This shows how the intervention, designed to guide a single targeted agent towards an additional desired outcome, is integrated into that agent’s decision-making process within the MAID framework. We present these implementation-specific MAIDs and their corresponding relevance graphs for both one-shot and sequential decision-making scenarios.
F.1 Pre-Strategy Intervention Implementation in One-Shot Decision Making
Figure 7 illustrates the MAID structures representing our implemented Pre-Strategy Intervention for one-shot decision-making in both simultaneous and sequential moves. These diagrams illustrate how the pre-decision node ($D_{pre}$) and the guidance signal/additional desired outcome variable ($Z$) (represented as an intrinsic reward) directly inform the targeted agent’s decision ($D_{h}$). Furthermore, an edge from a variable representing the additional desired outcome $Z$ to agent’s utility node $U_{h}$ indicates that the agent’s utility now explicitly incorporates this additional desired outcome.
F.2 Pre-Strategy Intervention Implementation in Sequential Decision Making
Figure 8(a) illustrates the MAID structures and its implementation for applying Pre-Strategy Intervention for sequential decision making with simultaneous moves. Similarly, Figure 8(b) shows these for sequential-move scenarios. These diagrams show the temporal extension, where the intervention influences decisions at each relevant timestep.
Appendix G Representing Markov Games through Multi-Agent Influence Diagrams
This appendix details how a standard team reward Markov game can be equivalently represented as a Multi-Agent Influence Diagram (MAID).
Markov game [26] is a popular mathematical model to describe the multi-agent sequential decision process across various real-world applications [84, 85, 86]. For succinct description, we only consider the team reward Markov game with a finite episode length, as described in Definition G.1.
Definition G.1 (Markov Game [35]).
A team reward Markov game can be described as a tuple $\langle\mathcal{I},\mathcal{S},\mathcal{A},T,R,L\rangle$. $\mathcal{I}$ is a set of agents; $\mathcal{S}$ is a set of states; $\mathcal{A}=\times_{i\in\mathcal{I}}\mathcal{A}_{i}$ is a set of joint actions and $\mathcal{A}_{i}$ is agent $i$’s action set; $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ describes the transition function that maps a state $s_{t}\in\mathcal{S}$ at timestep $t$ to $s_{t+1}\in\mathcal{S}$ at timestep $t+1$; $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is a team reward function evaluating the immediate joint action $a_{t}\in\mathcal{A}$ at some state $s_{t}\in\mathcal{S}$. In a team reward Markov game with an episode length of $L$ timesteps, agents aim to learn a joint policy $\pi=(\pi_{i})_{i\in\mathcal{I}}$ where $\pi_{i}:\mathcal{S}\rightarrow\mathcal{A}_{i}$ is agent $i$’s stationary policy, to solve the following optimization problem such that $\max_{\pi}\mathbb{E}_{\pi,T}[\sum_{t=0}^{L}R(s_{t},a_{t})]$.
A team reward Markov game can be rendered as a MAID, because we can match variables between these two models. In more details, both models’ agent sets are $\mathcal{I}$; $\mathcal{S}$ is associated with chance variables $\mathcal{X}$; $\mathcal{A}$ is associated with decision variables $\mathcal{D}$; $T$ is associated with conditional probability distributions $Pr$; $\pi$ is associated with decision rules $\delta$; and $\mathbb{E}_{\pi,Pr}[\sum_{t=0}^{L}R(s_{t},a_{t})]$ is associated with the expected utility shown in Equation 1. The objective of a team reward Markov game $\max_{\pi}\mathbb{E}_{\pi,Pr}[\sum_{t=0}^{L}R(s_{t},a_{t})]$ is equivalent to reaching a Nash equilibrium (see Definition 2.2), given that agents share total utility variables as team rewards.
Appendix H Implementation Details
The implementation is mainly based on JaxMARL [87]. Our experiments were run on NVIDIA RTX 4090 and A100 GPUs. A typical training run for the Hanabi environment takes approximately one hour on these systems.
This section details the architecture of our proposed method and the baselines used for comparison. Subsequently, we describe how the “additional desired outcome” is defined and implemented within the MPE and Hanabi environments respectively.
H.1 Implementation of Our Method
Our method, Pre-Strategy Intervention, is realized by augmenting a base MARL agent’s architecture with our novel pre-policy module. The full architecture is composed of three key components that work in concert: a standard network backbone, a Graph Neural Network (GNN) module for relational reasoning, and the pre-policy module for processing the intervention signal.
Agent Architecture. At each timestep, the agent’s raw observation is processed in parallel by two components. First, the GNN module constructs a graph representation of the observation to capture relational features between entities (e.g., agents, landmarks), which aligns with the principles of MAIDs (full implementation details are provided in Appendix H.3). Second, the pre-policy module takes the same observation along with the intrinsic reward from the previous step to produce a guidance embedding. Finally, the outputs from the standard backbone, the GNN embedding, and the pre-policy guidance embedding are concatenated and fed into the agent’s decision-making head (e.g., to compute Q-values or inform the critic).
This entire architecture, including the agent’s policy, the GNN module and the pre-policy module, is trained jointly end-to-end to maximize a composite shaping reward. The composite shaping reward is the sum of the extrinsic task reward and the intrinsic reward derived from the additional desired outcome. As established in Section 4.3, this training process is theoretically sound, as maximizing the intervention’s causal effect requires observing the total team utility ($U_{tot}$), which directly aligns with the use of a shared team reward in many previous MARL works.
Intrinsic Reward Function. The intrinsic reward quantifies the agent’s adherence to the specified “additional desired outcome” at each timestep. For example, in MPE this is the negative distance to a target landmark, while in Hanabi it is a value indicating compliance with a convention like “5 Save”. This intrinsic reward serves two purposes: (1) it is a direct input to the pre-policy module, and (2) it is added to the extrinsic task reward to form a composite shaping reward that guides the learning of the entire system.
Architectural Choice (Non-Parameter Sharing). Since our method focuses on the effect of intervening on a single targeted agent, other agents’ parameters update should be isolated from the targeted agent. For this reason, it is most naturally implemented with a non-parameter sharing (NP) architecture. The Pre-Strategy Intervention is designed to influence only the targeted agent, and an NP setup allows for this specialized learning without forcing homogeneity across all agents’ policies.
Target Agent Selection. For the experiments presented in this paper, we manually select a targeted agent in each environment. This experimental design allows for a clear and controlled analysis of the mechanism’s impact and mirrors practical scenarios where a human operator might manually choose a specific agent to intervene on.
H.2 Implementation of Baselines
Base MARL. This refers to the standard, unmodified (“vanilla”) version of the underlying MARL algorithm (e.g., IQL, VDN, IPPO). It uses the same network architecture and hyperparameters as our method but excludes both the pre-policy module and the additional intrinsic reward to maximize.
Intrinsic Reward. This approach serves as a direct ablation of our method to assess the contribution of the pre-policy module. In this variant, the pre-policy module is removed, while the additional desired outcome is conveyed to the single targeted agent implicitly via maximizing an intrinsic reward along with maximizing the extrinsic task reward.
Global Intervention. This approach represents a category of algorithms implementing the global intervention paradigm. In our experiments, we follow the specific implementation of two representative methods for this approach: LIIR [8] for policy-based algorithms, and LAIES [29] for value-based algorithms. Both methods in principle allocate an additional intrinsic reward to all agents to maximize. This contrasts the Intrinsic Reward approach above, where only one targeted agent is provided with an intrinsic reward to maximize.
H.3 GNN Implementation
Algorithm 1 Graph Embedding using GNNs
procedure GraphEmbedding(Observation)
  Input: Observation vector
  Represent the observation as an influence diagram with semantic features
  Apply graph convolution using a GNN with an adjacency matrix (either learned or manually crafted)
  Return: Graph embedding vector
end procedure
The foundation of our targeted intervention method (Pre-Strategy Intervention) is formed by the s-reachability graph criterion, introduced to identify the relevant policies of other agents for decision-making [16]. However, considering only policies is insufficient, as an agent’s policy may depend on other elements within the game [79]. Algorithm 1 provides an implementation how we can build connection with causal graph structure for pre-policy learning. By leveraging this graph structure, we incorporate prior knowledge about the game to help guide agents’ policy-making in practical. The feasibility of learning the causal graph during training has been demonstrated by Richens and Everitt [88], where agents can learn the causal model implicitly during interaction with the environment.
Architecture. If the adjacency matrix is not predefined, the GNN processes the observation vectors by first encoding them into logits, which are used to generate a soft adjacency matrix via the gumbel-softmax technique [89]. This matrix defines the relationship between features in the observations. Once the adjacency matrix is formed, a graph convolutional layer applies message passing to update the features of each node based on its neighbors [83]. The output node features are then aggregated using a mean-pooling operation to produce a graph embedding. This embedding is used for further processing or decision-making.
H.4 MPE
Multi-Agent Particle Environment (MPE). In the MPE Simple Spread task [18], our setup features 3 agents in a 2D continuous space cooperatively navigating to 3 landmarks while avoiding collisions. We apply our targeted intervention to one agent, using an intrinsic reward defined by the negative distance to its selected landmark. This is a simultaneous-move environment.
H.4.1 Intrinsic reward
In the environment, we define two scenarios for providing intrinsic rewards to the targeted agent.
Scenario 1 (Fixed Target Landmark). The targeted agent receives an intrinsic reward for approaching a specific, predetermined landmark. This is implemented by treating one landmark (e.g., the first one listed in the observation, regardless of its random initial position) as the target. The agent receives an additional reward inversely proportional to its distance to this fixed target landmark.
Scenario 2 (Dynamic Target Landmark - Farthest from Teammates). The targeted agent receives an intrinsic reward for approaching the landmark currently farthest from its other two teammates. This encourages strategic positioning relative to teammates. The farthest landmark $\mathbf{l}^{*}$ is determined dynamically at each step as follows:
Definition of the Farthest Landmark. For each landmark $\mathbf{l}\in\mathcal{L}$, compute the Euclidean distances to teammates $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$:
$d_{i}(\mathbf{l})=\|\mathbf{l}-\mathbf{p}_{i}\|_{2},\quad\text{for }i=1,2.$
Next, determine the minimum distance for each landmark:
$d_{\text{min}}(\mathbf{l})=\min\{d_{1}(\mathbf{l}),d_{2}(\mathbf{l})\}.$
Identify the farthest landmark $\mathbf{l}^{*}$ with the maximum minimum distance:
$\mathbf{l}^{*}=\arg\max_{\mathbf{l}\in\mathcal{L}}\,d_{\text{min}}(\mathbf{l}).$
The intrinsic reward for the targeted agent is then calculated as:
$\text{additional\_rew}=-\|\mathbf{p}_{0}-\mathbf{l}^{*}\|_{2},$
where $\mathbf{p}_{0}$ is the targeted agent’s position.
H.5 Hanabi
Hanabi. Team success in the Hanabi card game heavily relies on shared conventions—implicit rules that enhance predictability and coordination [19]. Our experiments study the 2-agent version of the game. We investigate if our targeted intervention method (Pre-Strategy Intervention), applied to a single agent, can effectively promote adherence to such conventions and thereby improve overall team coordination. Specifically, we guide the targeted agent towards one of two common Hanabi conventions as a additional desired outcome: “5 Save” (prioritizing hints for unique, high-value cards to prevent accidental discards) or “The Chop” (making discards predictable by targeting the newest, unhinted card).
H.5.1 Hanabi Environment and Experimental Setup Details
Hanabi is played by a team of 2-5 players (agents). The standard deck consists of cards in five different suits (colors), with cards in each suit numbered 1 to 5. There are three “1”s, two each of “2”s, “3”s, and “4”s, and one “5” for each suit, totaling 50 cards. The objective is for the team to collaboratively play cards in ascending order (1 through 5) for each of the five suits onto shared “firework” piles. Players hold a hand of cards (typically 4 or 5 cards, depending on the number of players), but crucially, they cannot see the cards in their own hand; they can only see the hands of their teammates.
On their turn, an agent must perform one of three actions:
  • Give a Hint: The agent can give a hint to a teammate about the cards in that teammate’s hand. To do so, they expend one of the team’s shared “hint tokens”. If no hint tokens are available, this action cannot be performed.
  • Discard a Card: The agent selects a card from their own (unknown) hand to discard to a common discard pile. This action recover one hint token. The discarded card is revealed to all players and is then out of play. After discarding, the agent draws a new card from the deck.
  • Play a Card: The agent selects a card from their own hand and attempts to play it onto one of the firework piles. If the card is playable (i.e., it is a “1” of a suit for which no card has been played, or it is the next number in sequence for a suit that has already been started), it is successfully added to the corresponding firework. If the card is not playable, it is discarded, and the team loses one of shared “lives”. If all lives are lost, the game ends immediately, and the team scores 0. After playing (successfully or not), the agent draws a new card.
The team’s score is the sum of the highest card values successfully played for each of the five suits.
The main difficulty in the Hanabi card game comes from the fact that players cannot see their own cards, a major limitation on what they know. Also, communication between players is very restricted and has a cost, as hints are limited. This requires players to make complicated deductions and makes working together effectively as a team extremely hard. For example, players need to discard cards to get more hint tokens. However, this is risky because they might accidentally remove unique and essential cards from the game, which could lower the team’s possible score.
Consequently, high-performing Hanabi teams, both human and AI, heavily rely on shared conventions. These are implicit behavioural rules or patterns that enhance the predictability and informativeness of actions, particularly for hints and discards, beyond their literal meaning. However, establishing consistent and effective conventions across all agents is a known and significant hurdle for standard multi-agent reinforcement learning (MARL) approaches, which may struggle to ensure all agents converge to and adhere to the same implicit rules.
H.5.2 Intrinsic Reward
In the environment, we implement two conventions for defining intrinsic rewards.
Convention 1 (5 Save): This convention is crucial for maximizing potential scores, as rank 5 cards are unique for each suit and essential for completing a firework. It aims to enhance team coordination by guiding agents to provide timely and clear hints about rank 5 cards held by teammates. This is especially important if those cards are currently unhinted and therefore at high risk of being accidentally discarded by a teammate unaware of their value. The objective is to ensure these vital cards are identified, preserved, and eventually played successfully. The specific logic for calculating the intrinsic reward, which encourages the targeted agent’s convention-adherent hinting actions (e.g., when a teammate holds an unhinted rank 5 card), is detailed in Algorithm 2.
Convention 2 (The Chop:) This convention promotes predictability in discarding actions, which in turn helps improve team efficiency and reduce miscoordination. It establishes a clear default discard target, often targeted as the newest card in a player’s hand that has not yet received any hint (the “chop” card). By adhering to this rule, especially when no other plays or hints are clearly beneficial, an agent provides a predictable pattern for discards. This allows teammates to make more informed inferences about the game state and their own hands. The intrinsic reward encourages the targeted agent to follow this specific discard protocol. The detailed logic for identifying the “chop” card and calculating the reward based on the agent’s discard action is provided in Algorithm 3.
Algorithm 2 “5 Save” Convention
1:procedure Calculate 5 Save Reward($a_{t}$, $T$, hands, hints)
2:  Input: action $a_{t}$, set of teammates $T$, hands hands, hint info hints.
3:  Output: Intrinsic reward $r_{5save}$.
4:  $r_{5save}\leftarrow 0$
5:  if any teammate $p\in T$ has an unhinted rank 5 card in hand then
6:   if $a_{t}$ is a hint action and $a_{t}$ does NOT hint rank “5” or the color of an unhinted 5 then
7:     $r_{5save}\leftarrow-1$ $\triangleright$ Penalty: Failed to prioritize hinting the 5
8:   end if
9:  end if
10:  Return: $r_{5save}$
11:end procedure
Algorithm 3 “The Chop” Convention
1:procedure Calculate The Chop Reward($a_{t},\text{hand},\text{hints}$)
2:  Input: current action $a_{t}$, agent’s hand hand, hint info hints.
3:  Output: Intrinsic reward $r_{chop}$.
4:  $r_{chop}\leftarrow 0$
5:  if $a_{t}$ is a discard action then
6:   Identify chop_card_indexchop\_card\_index using hand and hints (rightmost unhinted card index, or None if all hinted).
7:   Let discarded_indexdiscarded\_index be the index discarded by $a_{t}$.
8:   if chop_card_indexchop\_card\_index is None then
9:     $r_{chop}\leftarrow-1$ $\triangleright$ Discouraged: Discarding when all cards are hinted
10:   else if $discarded\_index\neq chop\_card\_index$ then
11:     $r_{chop}\leftarrow-2$ $\triangleright$ Penalty: Discarded a non-chop card
[…]
13:  end if
14:  Return: $r_{chop}$
15:end procedure
Appendix I Additional Experiments
I.1 MPE
I.1.1 Detailed Results of the Scenario shown in the Main Paper
We now present the detailed results for the MPE scenario discussed in the main paper, comparing our targeted intervention approach (Pre-Strategy Intervention) against base MARL methods. These comparisons are conducted across various MARL algorithms, with results shown in Figure 9. The baseline methods are implemented in two common settings: parameter sharing (PS) and non-parameter sharing (NP). Our Pre-Strategy Intervention is implemented using a non-parameter sharing architecture, reflecting more realistic scenarios where agents may not be able to directly exchange parameters during training.
(a) IQL (extrinsic).
(b) IQL (intrinsic).
(c) VDN (extrinsic).
(d) VDN (intrinsic).
(e) QMIX (extrinsic).
(f) QMIX (intrinsic).
Figure 9: Detailed results of the scenario of MPE shown in the main body of paper. For each base MARL algorithm (IQL, VDN, and QMIX), extrinsic and intrinsic returns are displayed side by side. The algorithm implemented in parameter sharing is denoted as PS, while that implemented in non-parameter sharing is denoted as NP.
I.1.2 Detailed Results of an Additional Scenario
In this scenario, the targeted agent receive additional feedback in intrinsic rewards, when it is close to the landmarks that are farthest from the another two teammates, as detailed in Appendix H.4. Figure 10 presents a detailed comparison of our targeted intervention approach (Pre-Strategy Intervention) against baseline methods, evaluated across various underlying MARL algorithms. Similarly, we consider baselines implemented in two popular implementation settings: parameter sharing (PS) and non-parameter sharing (NP). Our method, using non-parameter sharing, is observed to generally outperform the NP baselines by a significant margin in both primary task completion (extrinsic return) and attainment of the additional desired outcome (intrinsic return). Even compared with baselines implemented in PS, our method can still manifest comparable performance.
(a) IQL (extrinsic).
(b) IQL (intrinsic).
(c) VDN (extrinsic).
(d) VDN (intrinsic).
(e) QMIX (extrinsic).
(f) QMIX (intrinsic).
Figure 10: Detailed results of an additional scenario of MPE. For each base MARL algorithm (IQL, VDN, and QMIX), extrinsic and intrinsic returns are displayed side by side. The algorithm implemented in parameter sharing is denoted as PS, while that implemented in non-parameter sharing is denoted as NP. Our method is implemented in non-parameter sharing.
I.2 Hanabi
I.2.1 Detailed Results of the Scenario shown in the Main Paper
We now present detailed results for the Hanabi scenario discussed in the main paper (using the “5 Save” convention as the additional desired outcome), comparing our targeted intervention approach (Pre-Strategy Intervention) against base MARL methods. These comparisons are conducted across various underlying MARL algorithms, with results shown in Figure 11. The baseline methods are implemented in two common settings: parameter sharing (PS) and non-parameter sharing (NP). Our Pre-Strategy Intervention is implemented using a non-parameter sharing architecture, reflecting more realistic scenarios where agents may operate with decentralized learning and not directly exchange parameters during training.
(a) IPPO (extrinsic).
(b) IPPO (intrinsic).
(c) MAPPO (extrinsic).
(d) MAPPO (intrinsic).
(e) PQN-IQL (extrinsic).
(f) PQN-IQL (intrinsic).
(g) PQN-VDN (extrinsic).
(h) PQN-VDN (intrinsic).
Figure 11: Detailed results of the scenario of Hanabi shown in the main body of paper (with the convention as “5 Save”). For each base MARL algorithm (IPPO, MAPPO, PQN-IQL and PQN-VDN), extrinsic and intrinsic returns are displayed side by side. The algorithm implemented in parameter sharing is denoted as PS, while that implemented in non-parameter sharing is denoted as NP. Our method is implemented in non-parameter sharing.
I.2.2 Detailed Results of an Additional Scenario
We now consider the scenario with a new convention called “The Chop.” Similar to above, we present detailed results comparing our targeted intervention approach (Pre-Strategy Intervention) against baselines, across various underlying MARL algorithms, in Figure 12. The baselines again include implementations with parameter sharing (PS) and non-parameter sharing (NP). The results indicate that our method generally outperforms the non-parameter sharing (NP) baselines by a significant margin in both primary task completion (extrinsic return) and adherence to “The Chop” convention (intrinsic return). An exception is observed with the IPPO backbone, where the performance difference is less pronounced. Compared to parameter sharing (PS) baselines, our Pre-Strategy Intervention often demonstrates comparable performance.
(a) IPPO (extrinsic).
(b) IPPO (intrinsic).
(c) MAPPO (extrinsic).
(d) MAPPO (intrinsic).
(e) PQN-IQL (extrinsic).
(f) PQN-IQL (intrinsic).
(g) PQN-VDN (extrinsic).
(h) PQN-VDN (intrinsic).
Figure 12: Detailed results of the scenario of Hanabi shown in the main body of paper (with the convention as “The Chop”). For each base MARL algorithm (IPPO, MAPPO, PQN-IQL and PQN-VDN), extrinsic and intrinsic returns are displayed side by side. The algorithm implemented in parameter sharing is denoted as PS, while that implemented in non-parameter sharing is denoted as NP. Our method is implemented in non-parameter sharing.
I.3 Additional Results Analysis
Generalizability Across Different Additional Desired Outcomes and Base MARL Algorithms. As shown in Figures 9, 10, 11, and 12, our targeted intervention approach (Pre-Strategy Intervention) consistently outperforms baseline approaches across various base MARL algorithms. Additionally, the results across different experimental scenarios (each with distinct additional desired outcomes, such as specific landmark targeting in MPE or convention adherence in Hanabi) suggest that incorporating well-defined additional desired outcomes via the Pre-Strategy Intervention can effectively promote the agent learning process towards improved overall task performance.
Relationship Between Primary Task Goals and additional desired outcomes. We find that guiding the targeted agent (via the additional desired outcome) to move towards the landmark farthest from the other two teammates aligns with effective task completion in the MPE environment. As shown in Figure 10, the baselines exhibit an increasing trend in intrinsic reward compared to Figure 9, suggesting that moving towards the farthest landmark may share some overlap with the solution to the optimality of the task goal in MPE. However, in scenarios where the targeted agent’s additional desired outcome is a randomly pre-assigned (fixed) landmark, the baselines do not show a similar spontaneous increase in intrinsic reward for that specific fixed landmark, as observed in Figure 9. This may be because moving to the farthest landmark is not inherently required to accomplish the primary task goal. Instead, the predictable movement pattern of the targeted agent consistently favouring certain landmarks helps coordinate the team’s movement, implicitly facilitating task completion.
I.4 Additional Comparison with Global Intervention Paradigm in Hanabi
Figure 13: Comparison of extrinsic returns in Hanabi, evaluating our Pre-Strategy Intervention (PSI) implementing the targeted intervention paradigm against an approach implementing the global intervention paradigm (i.e. applying intrinsic rewards to all agents [29]) across various MARL algorithm backbones.
Figure 13 provides further evidence in the Hanabi environment, specifically utilizing value-based MARL algorithm backbones (PQN-VDN and PQN-IQL), for the comparison between our Pre-Strategy Intervention implementing the targeted intervention paradigm and an approach implementing the global intervention paradigm. The results for extrinsic returns (task completion) are consistent with the findings presented in the main paper (e.g., Figure 5(d), which used different backbones). Our targeted intervention approach, when applied with either PQN-VDN or PQN-IQL, generally leads to higher final performance compared to the chosen global intervention approach (LAIES [29]) where all agents are simultaneously guided.
I.5 Experiment with Noisy Observation
To evaluate the robustness of our targeted intervention approach (Pre-Strategy Intervention) under imperfect information that is common in many real-world MARL settings, we introduce noise into the agents’ observations. Specifically, we perturb the belief distributions agents form about cards in Hanabi, simulating scenarios with sensor noise or imperfect state estimation.
The experiment is conducted in Hanabi, using “The Chop” convention as the additional desired outcome guided by the intervention. For conciseness, we evaluate performance using the PQN-VDN backbone. We test the following three noise scenarios to assess robustness under different train-test conditions:
(1) Noise during Training and Testing. This simulates deploying the system in an environment with persistent observation noise, where the agent can potentially adapt during training.
(2) Noise during Training Only. This simulates training under noisy conditions but testing in a clearer environment. This tests robustness to a decrease in noise level compared to training conditions.
(3) Noise during Testing Only. This simulates training in relatively clean conditions but deploying in a noisy environment. This tests robustness to unexpected or higher levels of observation noise not seen during training.
(a) Noise for both training and test (extrinsic).
(b) Noise for both training and test (intrinsic).
(c) Noise for only training (extrinsic).
(d) Noise for only training (intrinsic).
(e) Noise for only test (extrinsic).
(f) Noise for only test (intrinsic).
Figure 14: Comparison of performance across scenarios with different noise conditions imposed on training and testing. The first row shows results when both training and testing environments contain noise, the second row shows results when only the training environment contains noise, and the third row shows results when only the testing environment contains noise. For each scenario, extrinsic and intrinsic returns are displayed side by side. The algorithm implemented in parameter sharing is denoted as PS, while that implemented in non-parameter sharing is denoted as NP. Our method is implemented in non-parameter sharing. The base MARL algorithm we select is PQN-VDN.
I.5.1 Both training and testing contain noise
In this scenario, the uniform noise (scaled by 0.2) perturbs the card belief distributions during both training and testing, simulating persistent observation uncertainty. As shown in Figures 14(a) and 14(b), our method still generally outperforms the base MARL algorithms in primary task completion (extrinsic return). Furthermore, the adherence to the additional desired outcome (intrinsic return) remains effective, indicating robustness when agents can adapt to consistent noise during training.
I.5.2 Only training environments contain noise
In this scenario, the noise (scaled by 0.05) is present only during training, simulating a scenario where the deployment environment has less noise than the training conditions. Figures 14(c) and 14(d) show that our method maintains an advantage in task completion over the baselines. However, the performance regarding the additional desired outcome (intrinsic return) is notably impaired compared to the noise-free or consistently noisy cases, performing similarly to the non-parameter sharing baseline.
I.5.3 Only testing environments contain noise
This scenario introduces the noise (scaled by 0.02) only during testing, simulating deployment in an environment with unexpected observation noise. As seen in Figures 14(e) and 14(f), the performance of our method aligns closely with the parameter-sharing baseline in terms of extrinsic return (with overlapping confidence intervals). Similarly, there is no significant advantage in achieving the additional desired outcome (intrinsic return) compared to the non-parameter sharing baseline.
I.6 Hanabi 4 Players
(a) 4-player Hanabi (extrinsic).
(b) 4-player Hanabi (intrinsic).
Figure 15: Performance comparison in 4-player Hanabi, with PQN-VDN as the base algorithm for all methods. Our Pre-Strategy Intervention (PSI) is compared against the Base MARL, the Global Intervention and the Intrinsic Reward baselines.
To further validate the scalability of our Pre-Strategy Intervention, we conducted new experiments on the 4-player version of the Hanabi challenge. This setting serves as a more demanding benchmark for coordination, as the strategic complexity and communication burden are known to increase dramatically with the number of players [90]. The results presented in Figure 15 demonstrate that our method achieves significantly higher scores than both the base MARL algorithms and the approach of the global intervention paradigm. This strong performance in a larger-scale setting provides robust evidence of our method’s ability to effectively handle more complex coordination challenges.
I.7 Heterogeneous MPE
(a) Heterogeneous MPE (extrinsic).
(b) Heterogeneous MPE (intrinsic).
Figure 16: Performance evaluation in a Heterogeneous MPE scenario. We compare the effect of applying our Pre-Strategy Intervention to a fast agent versus a normal-speed agent, with IQL as the base algorithm, against a baseline without intervention (Base MARL). The heterogeneous environment consists of one fast and two normal-speed agents.
To evaluate the performance of our targeted intervention approach (Pre-Strategy Intervention) in settings with diverse agent capabilities, we designed a heterogeneous MPE scenario. This environment features three agents with asymmetric speeds: one “sprinter” agent with five times the normal speed and two normal-speed agents, with all methods built upon the IQL algorithm. The results, shown in Figure 16, reveal a consistent trend: intervening on a normal-speed agent leads to a noticeable improvement in final team performance (extrinsic return) compared to intervening on the uniquely capable sprinter, although the magnitude of this difference is modest in this specific environment. Our analysis suggests this is because the system’s primary limitation: the coordination bottleneck between the two normal-speed agents, when intervening on the sprinter agent. Even a subtle increase in the predictability of one normal agent’s behavior through our intervention can alleviate this core challenge, leading to a more effective overall team strategy and demonstrating our approach’s potential in heterogeneous teams where targeted coordination adjustments can yield benefits.
The above results and analysis suggest that further investigation is needed to determine which agent to intervene on in future work.
Appendix J Broader Impacts
Our work on Pre-Strategy Intervention (implementing the targeted intervention paradigm) in multi-agent reinforcement learning (MARL) introduces a theoretical framework for agent coordination. As foundational research conducted entirely in simulated environments, the immediate societal impact is limited. However, we acknowledge that theoretical advances often lay groundwork for future applications. If extended to practical domains, this research could potentially enable more efficient coordination in multi-agent systems by focusing guidance on specific agents rather than requiring direct control of an entire system. Our targeted intervention paradigm could reduce computational and communication overhead in domains like robotic teams, traffic management, or resource allocation. At the same time, we recognize that techniques for influencing multi-agent systems through targeted intervention could potentially be misapplied if deployed in socially sensitive contexts without appropriate oversight. Future work that moves this theoretical foundation toward real-world applications should carefully consider application-specific ethical implications and incorporate appropriate safeguards against potential misuse. As our current work remains at the algorithmic level with experiments limited to standard MARL benchmarks (MPE and Hanabi), we do not anticipate immediate societal consequences from this research contribution.
Appendix K Hyperparameters
We show the hyperparameters for all experiments we conduct for the ease of reproducing results.
Table 2: Hyperparameters for MPE in IQL across two scenarios.
Hyperparameter Scenario 1 Scenario 2
Total Timesteps $2\times 10^{6}$ $2\times 10^{6}$
Number of Environments 8 8
Number of Steps 26 26
Buffer Size 5000 5000
Buffer Batch Size 32 32
Network Hidden Size 64 64
Epsilon Start 0.8 0.6
Epsilon Finish 0.02 0.1
Epsilon Decay 0.1 0.1
Learning Rate (LR) 0.0035 0.0045
Learning Starts 10000 timesteps 10000 timesteps
Gamma ($\gamma$) 0.9 0.9
Max Gradient Norm 5 5
Target Update Interval 200 200
Tau ($\tau$) 1.0 1.0
Number of Epochs 1 1
Intrinsic Reward Ratio 0.6 0.2
Pre-Policy Module Output Size 16 32
Pre-Policy Module Hidden Size 64 64
GNN Node Embedding Size 16 16
Table 3: Hyperparameters for MPE in VDN across two scenarios.
Hyperparameter Scenario 1 Scenario 2
Total Timesteps $2\times 10^{6}$ $2\times 10^{6}$
Number of Environments 8 8
Number of Steps 26 26
Buffer Size 5000 5000
Buffer Batch Size 32 32
Network Hidden Size 64 64
Epsilon Start 0.8 0.6
Epsilon Finish 0.1 0.1
Epsilon Decay 0.1 0.1
Learning Rate (LR) 0.007 0.0045
Learning Starts 10000 timesteps 10000 timesteps
Gamma ($\gamma$) 0.9 0.9
Max Gradient Norm 5 5
Target Update Interval 200 200
Tau ($\tau$) 1.0 1.0
Number of Epochs 1 1
Intrinsic Reward Ratio 0.6 0.2
Pre-Policy Module Output Size 32 32
Pre-Policy Module Hidden Size 64 64
GNN Node Embedding Size 16 16
Table 4: Hyperparameters for MPE in QMIX.
Hyperparameter Value
Total Timesteps $2\times 10^{6}$
Number of Environments 8
Number of Steps 26
Buffer Size 5000
Buffer Batch Size 32
Network Hidden Size 64
Epsilon Start 0.6
Epsilon Finish 0.1
Epsilon Decay 0.1
Learning Rate (LR) 0.004
Learning Starts 10000 timesteps
Gamma ($\gamma$) 0.9
Max Gradient Norm 5
Target Update Interval 200
Tau ($\tau$) 1.0
Number of Epochs 1
Intrinsic Reward Ratio 0.2
Pre-Policy Module Output Size 32
Pre-Policy Module Hidden Size 64
GNN Node Embedding Size 8
Mixer Embedding Size 32
Mixer Hypernetwork Hidden Size 128
Mixer Initialization Scale 0.001
Table 5: Hyperparameters for Independent PPO in MPE.
Hyperparameter Value
Total Timesteps $1\times 10^{7}$
Number of Environments 16
Number of Steps 128
Learning Rate 0.00055
Hidden Size 64
Update Epochs 4
Number of Minibatches 8
Gamma 0.99
GAE Lambda 0.85
Clip Epsilon 0.2
Scale Clip Epsilon False
Entropy Coefficient 0.015
Value Function Coefficient 0.95
Max Gradient Norm 10
Activation Tanh
Learning Rate Annealing True
Intrinsic Reward Ratio 0.2
Pre-Policy Module Hidden Size 64
Pre-Policy Module Output Size 16
GNN Node Embedding Size 8
Table 6: Hyperparameters for PQN in Hanabi.
Hyperparameter Value
Total Timesteps $1\times 10^{9}$
Number of Environments 1024
Number of Steps 1
Hidden Dimension 512
Number of Layers 3
Normalization Type Layer Normalization
Normalize Inputs False
Dueling True
Epsilon Start 0.03
Epsilon Finish 0.005
Epsilon Decay 0.05
Max Gradient Norm 0.5
Number of Minibatches 1
Number of Epochs 1
Learning Rate 0.0002
Learning Rate Linear Decay True
Lambda 0.0
Gamma 0.99
Intrinsic Reward Ratio 0.02
Pre-Policy Module Hidden Size 512
Pre-Policy Module Output Size 128
GNN Node Embedding Size 8
GNN Observation Encoder Hidden Size 64
GNN Adjacency Matrix Sampling Temperature 1.0
Table 7: Hyperparameters for IPPO in Hanabi.
Hyperparameter Value
Total Timesteps $1\times 10^{9}$
Number of Environments 1024
Number of Steps 128
Learning Rate 0.0006
Anneal Learning Rate False
Update Epochs 8
Number of Minibatches 8
Gamma 0.99
GAE Lambda 0.9
Clipping Epsilon 0.1
Entropy Coefficient 0.02
Value Function Coefficient 0.5
Max Gradient Norm 0.5
Intrinsic Reward Ratio 0.035
Pre-Policy Module Hidden Size 256
Pre-Policy Module Output Size 128
GNN Node Embedding Size 8
GNN Observation Encoder Hidden Size 128
GNN Adjacency Matrix Sampling Temperature 1.0
Table 8: Hyperparameters for MAPPO in Hanabi.
Hyperparameter Value
Total Timesteps $1\times 10^{9}$
Number of Environments 1024
Number of Steps 128
Learning Rate 0.0065
Anneal Learning Rate True
Number of Update Epochs 4
Number of Minibatches 4
Gamma 0.99
GAE Lambda 0.95
Clipping Epsilon 0.1
Entropy Coefficient 0.01
Value Function Coefficient 0.5
Max Gradient Norm 0.5
Activation Function ReLU
Intrinsic Reward Ratio 0.035
Pre-Policy Module Hidden Size 128
Pre-Policy Module Output Size 256
GNN Node Embedding Size 8
GNN Observation Encoder Hidden Size 128
GNN Adjacency Matrix Sampling Temperature 1.0